What is an equation of the line with y-intercept -15 and is perpendicular to the line whose equation is y=1/3+6
Answer by math_helper(2461)

What is an equation of the line with y-intercept -15 and is perpendicular to the line whose equation is y=1/3+6
—————————————————
NOTE:
Assuming the equation of the given line was intended to be written
y = (1/3)x + 6 (there is no x in the original problem statement)
—————————————————
Slope-intercept form is: y = mx + b (m=slope, b= y-axis intercept)
We know straight away that b=-15 as the problem asks for that.
As for m, two lines are perpendicular when m1 = -1/m2. Since the given
line has slope (1/3), the slope we want for the perpendicular line is -1/(1/3) = -3.

Ans: y = -3x -15